I need help with this one, the 0 is confusing me.
Solve by completing the square:
X2 –12X + 27 = 0

Answer by venugopalramana(3286)

I need help with this one, the 0 is confusing me.
Solve by completing the square:
X2 –12X + 27 = 0
{X^2-2(X)(6)+6^2-6^2}+27=0
(X-6)^2-6^2+27=0
(X-6)^2=36-27=9
X-6=+3.................OR..............X-6=-3
X=9...........................OR.................X=3